strinas:

What is the value of the expression (–5) -3? 1. Apply the negative exponents rule:     1 (−5)3 2. Expand the power:           1 (−5)(−5)(−5) 3. Simplify:                1 x What is the value of x? x =

strinas:

plz help

AZ:

what is \( (-5) * (-5) * (-5) = ?\)

AZ:

Remember a negative number times negative number gives us a positive number and a positive number times a negative number gives us a negative number

AZ:

This summarizes it pretty well

1 attachment
strinas:

ok thanks for the help

AZ:

My pleasure!

Angle:

@AZ a tiny mistake here is that I think this is a negative exponent problem

Angle:

hmm I take that back, maybe I'm reading the formatting wrong good job either way tho

AZ:

I understood that there was a formatting error when copy-pasting the question and took that into consideration. This is how I understood the question to be: \((-5)^{-3} = \Large\frac{1}{(-5)^{3}} = \frac{1}{(-5)\times (-5) \times (-5)} = \frac{1}{x}\)

Angle:

yeah, that's my bad then you've got a good eye to be able help with the question despite the formatting

AZ:

haha, thank you!